OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5) A. y + 5 = -1(x – 2) B. y + 2 = 1(x – 5) C. y + 1 = 5 (x + 2) D. y – 5 = -1(x + 2)

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope (m) = -1 x1 = -2 and y1 = 5 now lets sub in our info y - 5 = -1(x - (-2) y = 5 = -1 (x + 2) <-- your answer

OpenStudy (texaschic101):

oops...y - 5 = -1(x + 2)

OpenStudy (anonymous):

Thanks for the help

OpenStudy (texaschic101):

anytime :)